# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 4.1 Absorption theorem
For simplicity, we sketch the proof for a slighty simplified version with $\mathbf{A}=\mathbf{C}$. So, we assume that $\mathbf{A}$ is Taylor and that $R\leq_{sd}\mathbf{A}^{2}$ is linked. The original proof that $\mathbf{A}$ then necessarily has a proper absorbing subuniverse can be divided into 3 steps.
@@
    From $\mathbf{A}$ being Taylor it is derived that $\mathbf{A}$ either has a nontrivial 2-absorbing subuniverse or a transitive term operation $t$ of some arity $n$, i.e., for each $b,c\in A$ and every coordinate $i$ of $t$, there exists a tuple ${\mathbf{a}}\in A^{n}$ with $a_{i}=b$ such that $t({\mathbf{a}})=c$.
@@
    Using the transitive operation, it is proved that if $\mathbf{A}$ has no nontrivial absorbing subuniverses, then $R$ is left or right central.
@@
    It is shown that the transitive operation witnesses that the left (right) center absorbs $\mathbf{A}$.
The first step was explored in more detail in [8]. Lemma 2.7 in [8] shows that each algebra has a nontrivial projective subuniverse or a transitive term operation. A simple argument then shows that every projective subuniverse in a Taylor algebra is 2-absorbing, a witness is, e.g., any operation of the form $t(x,\dots,x$, $y,\dots,y)$ where $t$ is cyclic.
@@
The third step, that a transitive operation witnesses absorption of left centers, is straightforward, see Proposition 8.3. A significant refinement, Corollary 7.10.2 in [60], shows that left centers are, in fact, 3-absorbing. An adjustment of the proof will also help us in proving Theorem 5.10.
@@
([60], link to proof) If $B$ is a Taylor center of an algebra $\mathbf{A}$, then $B\trianglelefteq_{3}\mathbf{A}$.
Note that in the previous proposition $\mathbf{A}$ need not be a Taylor algebra, but $\mathbf{C}$ (where the witnessing relation is $R\leq_{sd}\mathbf{A}\times\mathbf{C}$) must be. The following proposition states that we can switch the condition:
@@
(link to proof) If $B$ is a center of a Taylor algebra $\mathbf{A}$, then $B\trianglelefteq_{3}\mathbf{A}$.
The assumptions of the latter proposition are easier to satisfy — the algebra $\mathbf{A}$ is usually Taylor by default.
@@
(link to proof) Suppose $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ is proper and linked, and $\mathbf{C}$ is Taylor. Then $\mathbf{A}$ or $\mathbf{C}$ has a nontrivial 3-absorbing subuniverse.
@@ 4.2 Subdirect irredundant subpowers
We now present the unification result. It says that any “interesting” (subdirect irredundant proper) relation either pp-defines an interesting binary relation or pp-defines (and is even inter-pp-definable with) ternary relations of very particular shape (strongly functional). We have recently found out that Theorem 4.7 can be also deduced from Szendrei’s results in [58]. Our contribution is thus the presented formulation, its alternative proof, and the application to the fundamental theorems of the dichotomy proofs.
@@
Theorem 4.7 implies that every algebra $\mathbf{A}$ has at least one of the following properties of its invariant relations.
@@
    $\mathbf{A}$ has no proper irredundant subdirect subpowers.
@@
    $\mathbf{A}$ has a proper irredundant binary subdirect subpower.
@@
    $\mathbf{A}$ has a ternary strongly functional subpower.
In the last case, it is easy to pp-define a congruence on $\mathbf{A}^{2}$ such that the diagonal is one of its classes, so $\mathbf{A}$ is abelian in that case.
@@
(link to proof) If $R\leq\mathbf{A}^{3}$ is a strongly functional relation, then $\mathbf{A}$ is abelian.
In case (1), subdirect relations have a very simple structure; for instance, any constraint $R(x_{1},\dots$, $x_{n})$ with subdirect $R$ is effectively a conjunction of bijective dependencies $x_{i}=f(x_{j})$ (in particular, $\mathbf{A}$ is simple). It is also immediate that $\mathbf{A}$ is polynomially complete. Less trivially, case (1) often leads to majority edges, as we show in Theorem 4.10 below. We require the following definition first.
@@
Let $\mathbf{A}$ be an algebra. By the connected-by-subuniverses equivalence, denoted $\mu_{\mathbf{A}}$, we mean the smallest equivalence containing all the pairs $(a,b)$ such that $\operatorname{Sg}_{\mathbf{A}}(a,b)\neq A$.
We remark that the equivalence $\mu_{\mathbf{A}}$ is not, in general, a congruence of $\mathbf{A}$, so this concept may seem somewhat unnatural from the algebraic perspective. This equivalence may be full. If we are in case (1) and $\mu_{\mathbf{A}}$ is not full, then the following theorem produces majority edges.
@@
(link to proof) Suppose that $\mathbf{A}$ has no subdirect proper irredundant subpowers. Then there exists a term operation $t\in\operatorname{Clo}_{3}(\mathbf{A})$ such that for any $a,b\in A$ with $\operatorname{Sg}_{\mathbf{A}}(a,b)=\mathbf{A}$ (e.g., $(a,b)\notin\mu_{\mathbf{A}}$), $t(a,a,b)=t(a,b,a)=t(b,a,a)=a$.
In case (2) and when $\mathbf{A}$ is simple, a binary irredundant relation is necessarily linked. Then we get a central relation, e.g., by Proposition 4.2, and, by the following theorem, also semilattice edges in case that $\mu_{\mathbf{A}}$ is not full and the size of $A$ is at least three (in case that $\mu_{\mathbf{A}}$ is full the theorem is vacuously true).
@@
(link to proof) Suppose $\mathbf{A}$ with $|A|>2$ is simple and there exists a proper irredundant subdirect binary subpower. Then there exists a $\mu_{\mathbf{A}}$-class $B$ such that, for every $b\in B,a\notin B$, the pair $(a,b)$ is a semilattice edge witnessed by the identity congruence.
Zhuk’s four types theorem (Theorem 3.7) is now a consequence of Theorem 4.7, Proposition 4.8, and Proposition 4.2. Indeed, one simply applies these facts to $\mathbf{A}$ factored by a maximal congruence, which is a simple algebra, and then lifts 2-absorbing subuniverses and centers back to $\mathbf{A}$. Moreover, we can improve item (d) in the four types theorem as follows.
@@
(link to proof) Let $\mathbf{A}$ be an algebra, then
@@
    $\mathbf{A}$ has a nontrivial 2-absorbing subuniverse, or
@@
    $\mathbf{A}$ has a nontrivial center (which is a Taylor center in the case where $\mathbf{A}$ is a Taylor algebra), or
@@
    $\mathbf{A}/\alpha$ is abelian for some proper congruence $\alpha$ of $\mathbf{A}$, or
@@
    $\mathbf{A}/\alpha$ has no proper irredundant subdirect subpowers for some proper congruence $\alpha$ of $\mathbf{A}$.
@@ 5 Minimal Taylor algebras
An algebra $\mathbf{A}$ is called a minimal Taylor algebra if it is Taylor but no proper reduct of $\mathbf{A}$ is.
There are exactly four minimal Taylor algebras on a two-element set, up to term-equivalence: the two semilattices, the majority algebra, and the affine Mal’cev algebra. Indeed, from the description of their term operations given in Subsection 2.2 it follows that clones of these algebras are even minimal, in the sense that the only proper subclone is the clone of projections. A nice example of a minimal Taylor algebra on a three-element domain is the rock-paper-scissors algebra mentioned in Subsection 3.3. To see that this algebra is minimal Taylor observe that any term operation behaves on any two-element set like the term operation of a two-element semilattice with the same set of essential coordinates. Therefore, the original operation can be obtained by identifying variables in any term operation having at least two essential coordinates. The same argument shows that any semilattice, not necessarily two-element, is minimal Taylor. Affine Mal’cev algebras are also minimal Taylor; this can be deduced e.g. from Theorem 2.4.
In Subsection 5.1 we give the basic general theorems that were proved in [18] in the context of minimal bounded width algebras. Subsection 5.2 concentrates on absorption and related concepts in Zhuk’s theory. It turns out that 2-absorbing sets are exactly projective subuniverses (Theorem 5.7) and 3-absorbing sets are exactly centers (Theorem 5.10). Subsection 5.3 shows that edges substantially simplify in minimal Taylor algebras (Theorem 5.13) and gives additional information for minimal edges; in particular, minimal semilattice edges coincide with thin semilattice edges as defined in [26, 29] (Proposition 5.15). Finally, in Subsection 5.4, we demonstrate a strong interaction between absorption and edges. We show that 2-absorbing subuniverses are exactly subsets that are, in some sense, stable under all the edges (Theorem 5.19), we provide somewhat weaker interaction between absorbing subuniverses and subsets stable under semilattice and abelian edges (Theorem 5.21), we give a common witnessing operation for all of the edges as well as all of the 2- and 3-absorbing subuniverses (Theorem 5.23), and we show that each such a witnessing operation generates the whole clone of term operations (Theorem 5.24). Proofs for this section, including verification of examples, are in Section 9.
@@ 5.1 General facts
Another simple, but important consequence of cyclic operations is the following proposition. The result is slightly more technical than most of the others, but it is in the core of many strong properties of minimal Taylor algebras. The proposition can be further strengthened, see [19, Theorem 4.2.4.].
@@
(link to proof) Let $\mathbf{A}$ be a minimal Taylor algebra and $B\subseteq A$ be closed under an operation $f\in\operatorname{Clo}(\mathbf{A})$ such that $B$ together with the restriction of $f$ to $B$ forms a Taylor algebra. Then $B$ is a subuniverse of $\mathbf{A}$.
@@ 5.2 Absorption
(link to proof) Let $\mathbf{A}$ be a minimal Taylor algebra and $B$ an absorbing set of $\mathbf{A}$. Then $B$ is a subuniverse of $\mathbf{A}$.
@@
Let $\mathbf{A}$ be an algebra and $B\subseteq A$. The set $B$ is a strongly projective subuniverse of $\mathbf{A}$ if for every $f\in\operatorname{Clo}_{n}(\mathbf{A})$ and every essential coordinate $i$ of $f$, we have $f({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ is such that $a_{i}\in B$.
@@
The main value of this theorem is the implication showing that, in minimal Taylor algebras, every $2$-absorption, i.e. (a), is as strong as possible (d). Moreover (b) provides a nice relational description of $2$-absorption, which collapses the general condition from Proposition 2.6 for projectivity to arity $3$. The implications (d) implies (c) implies (b) implies (a) hold for all Taylor algebras. The following examples disprove the reverse implications. We leave the verification as an exercise.
@@
Let $t(x,y,z,w)$ be an operation on $\{0,1\}$ which is equal to $x\vee y\vee z\vee w$ everywhere, except that it is zero on the tuple $(1,0,0,0)$ and its cyclic shifts. Then $\mathbf{A}=(\{0,1\};t)$ is a Taylor algebra and the $R$ defined for $B=\{1\}$ in (b) of Theorem 5.7 is compatible with $\mathbf{A}$. On the other hand $B$ is not a projective subuniverse of $\mathbf{A}$ as witnessed by, for example, $t$ and thus (c) of Theorem 5.7 is false.
@@
    If $B\trianglelefteq_{2}\mathbf{A}$ and $C\leq\mathbf{A}$, then $B\cup C\leq\mathbf{A}$.
@@
    If $B\trianglelefteq_{2}\mathbf{A}$ and $C\trianglelefteq\mathbf{A}$ by $f$, where $\emptyset\neq B,C\neq A$, then
@@
    If $\mathbf{C}\trianglelefteq_{2}\mathbf{B}\trianglelefteq_{2}\mathbf{A}$, then $\mathbf{C}\trianglelefteq_{2}\mathbf{A}$.
@@
    $\mathbf{A}$ has a unique minimal 2-absorbing subalgebra $\mathbf{B}$. Moreover, this algebra $\mathbf{B}$ does not have any nontrivial 2-absorbing subalgebra.
@@
    there exists $\mathbf{C}$ with $\operatorname{Clo}(\mathbf{C})\subseteq\operatorname{Clo}(\{0,1\};\operatorname{maj})$ such that $R(x,y)=B(x)\vee(y=0)$ is a centrality witness.
@@
Just like in Theorem 5.7 we have that a relatively weak notion of $3$-absorption implies a very strong type of centrality which is (d). Let us investigate (d) in greater detail. The fact that $R$ is a subuniverse of $\mathbf{A}\times\mathbf{C}$ translates to the following fact. To every operation of $\mathbf{A}$, say $f$, we can associate an operation $f^{\prime}\in\operatorname{Clo}(\{0,1\},\operatorname{maj})$ such that $f({\mathbf{a}})\in B$ whenever $f^{\prime}({\mathbf{x}})=1$ and ${\mathbf{x}}$ is the characteristic tuple of ${\mathbf{a}}$ with respect to $B$ (i.e. $x_{i}=1$ if and only if $a_{i}\in B$). That is, from the viewpoint of “being outside $B$” vs. “being inside $B$” every operation outputs “inside $B$” every time the corresponding operation of $\operatorname{Clo}(\{0,1\};\operatorname{maj})$ outputs $1$.
In fact, there exists a cyclic $t$ in $\mathbf{A}$ (say, $p$-ary) such that $R(x,y)=B(x)\vee(y=0)$ is a subuniverse of $(A;t)\times(\{0,1\};\operatorname{maj}_{p})$ for every $3$-absorbing $B$. This translates to a simpler statement: for every $3$-absorbing $B$ we have $t({\mathbf{a}})\in B$ whenever majority of the $a_{i}$ belong to $B$, and we cannot expect more, as witnessed by the $2$-element majority algebra. Since $t$ is cyclic it generates the whole clone and, for example, item (2) in Proposition 5.12 below becomes obvious.
@@
(link to verification) Consider the algebra $\mathbf{A}=(\{0,1,2\},m)$ where $m$ is the majority operation such that $m(a,b,c)=a$ whenever |$|\{a,b,c\}|=3$. This algebra is minimal Taylor and the set $C=\{0,1\}$ is an absorbing subuniverse of $\mathbf{A}$. However, $C$ is not a center of $\mathbf{A}$.
@@
    If $B,C\trianglelefteq_{3}\mathbf{A}$, then $B\cup C\leq\mathbf{A}$ and $B\cap C\trianglelefteq_{3}\mathbf{A}$.
@@
    If $\emptyset\neq B,C\trianglelefteq_{3}\mathbf{A}$ and $B\cap C=\emptyset$, then $B^{2}\cup C^{2}$ is a congruence on the subalgebra of $\mathbf{A}$ with universe $B\cup C$ and the quotient is term-equivalent to a two-element majority algebra.
@@
    If $\mathbf{C}\trianglelefteq_{3}\mathbf{B}\trianglelefteq_{3}\mathbf{A}$, then $\mathbf{C}\trianglelefteq_{3}\mathbf{A}$.

### v6

@@ 4.1 Absorption theorem
For simplicity, we sketch the proof for a slighty simplified version with . So, we assume that is Taylor and that $R\leq_{sd}\mathbf{A}^{2}$ is linked. The original proof that then necessarily has a proper absorbing subuniverse can be divided into 3 steps.
@@
    From being Taylor it is derived that either has a nontrivial 2-absorbing subuniverse or a transitive term operation $t$ of some arity $n$, i.e., for each $b,c\in A$ and every coordinate $i$ of $t$, there exists a tuple ${\mathbf{a}}\in A^{n}$ with $a_{i}=b$ such that $t({\mathbf{a}})=c$.
@@
    Using the transitive operation, it is proved that if has no nontrivial absorbing subuniverses, then $R$ is left or right central.
@@
    It is shown that the transitive operation witnesses that the left (right) center absorbs .
The first step was explored in more detail in [Barto15:malcev]. Lemma 2.7 in [Barto15:malcev] shows that each algebra has a nontrivial projective subuniverse or a transitive term operation. A simple argument then shows that every projective subuniverse in a Taylor algebra is 2-absorbing, a witness is, e.g., any operation of the form $t(x,\dots,x$, $y,\dots,y)$ where $t$ is cyclic.
@@
The third step, that a transitive operation witnesses absorption of left centers, is straightforward, see Proposition 8.4. A significant refinement, Corollary 7.10.2 in [Zhuk20:dichotomy], shows that left centers are, in fact, 3-absorbing. An adjustment of the proof will also help us in proving Theorem 5.10.
@@
([Zhuk20:dichotomy], link to proof) If $B$ is a Taylor center of an algebra , then $B\trianglelefteq_{3}$.
Note that in the previous proposition need not be a Taylor algebra, but  (where the witnessing relation is $R\leq_{sd}$) must be. The following proposition states that we can switch the condition:
@@
(link to proof) If $B$ is a center of a Taylor algebra , then $B\trianglelefteq_{3}$.
The assumptions of the latter proposition are easier to satisfy — the algebra is usually Taylor by default.
@@
(link to proof) Suppose $R\leq_{sd}$ is proper and linked, and is Taylor. Then or has a nontrivial 3-absorbing subuniverse.
@@ 4.2 Subdirect irredundant subpowers
We now present the unification result. It says that any “interesting” (subdirect irredundant proper) relation either pp-defines an interesting binary relation or pp-defines (and is even inter-pp-definable with) ternary relations of very particular shape (strongly functional). We have recently found out that Theorem 4.7 can be also deduced from Szendrei’s results in [Szendrei:idempotent]. Our contribution is thus the presented formulation, its alternative proof, and the application to the fundamental theorems of the dichotomy proofs.
@@
Theorem 4.7 implies that every algebra has at least one of the following properties of its invariant relations.
@@
    has no proper irredundant subdirect subpowers.
@@
    has a proper irredundant binary subdirect subpower.
@@
    has a ternary strongly functional subpower.
In the last case, it is easy to pp-define a congruence on such that the diagonal is one of its classes, so is abelian in that case.
@@
(link to proof) If $R\leq$ is a strongly functional relation, then is abelian.
In case (1), subdirect relations have a very simple structure; for instance, any constraint $R(x_{1},\dots$, $x_{n})$ with subdirect $R$ is effectively a conjunction of bijective dependencies $x_{i}=f(x_{j})$ (in particular, is simple). It is also immediate that is polynomially complete. Less trivially, case (1) often leads to majority edges, as we show in Theorem 4.10 below. We require the following definition first.
@@
Let $\mathbf{A}$ be an algebra. By the connected-by-subuniverses equivalence, denoted $\mu_{\mathbf{A}}$, we mean the smallest equivalence containing all the pairs $(a,b)$ such that $\operatorname{Sg}(a,b)\neq A$.
We remark that the equivalence $\mu$ is not, in general, a congruence of , so this concept may seem somewhat unnatural from the algebraic perspective. This equivalence may be full. If we are in case (1) and $\mu$ is not full, then the following theorem produces majority edges.
@@
(link to proof) Suppose that has no subdirect proper irredundant subpowers. Then there exists a term operation $t\in\operatorname{Clo}_{3}($ such that for any $a,b\in A$ with $\operatorname{Sg}(a,b)=$ (e.g., $(a,b)\notin\mu_{\mathbf{A}}$), $t(a,a,b)=t(a,b,a)=t(b,a,a)=a$.
In case (2) and when is simple, a binary irredundant relation is necessarily linked. Then we get a central relation, e.g., by Proposition 4.2, and, by the following theorem, also semilattice edges in case that $\mu$ is not full and the size of $A$ is at least three (in case that $\mu$ is full the theorem is vacuously true).
@@
(link to proof) Suppose with $|A|>2$ is simple and there exists a proper irredundant subdirect binary subpower. Then there exists a $\mu$-class $B$ such that, for every $b\in B,a\notin B$, the pair $(a,b)$ is a semilattice edge witnessed by the identity congruence.
Zhuk’s four types theorem (Theorem 3.7) is now a consequence of Theorem 4.7, Proposition 4.8, and Proposition 4.2. Indeed, one simply applies these facts to factored by a maximal congruence, which is a simple algebra, and then lifts 2-absorbing subuniverses and centers back to . Moreover, we can improve item (d) in the four types theorem as follows.
@@
(link to proof) Let be an algebra, then
@@
    has a nontrivial 2-absorbing subuniverse, or
@@
    has a nontrivial center (which is a Taylor center in the case where is a Taylor algebra), or
@@
    is abelian for some proper congruence $\alpha$ of , or
@@
    has no proper irredundant subdirect subpowers for some proper congruence $\alpha$ of .
@@ 5 Minimal Taylor algebras
An algebra $\mathbf{A}$ is called a minimal Taylor algebra if it is Taylor but no proper reduct of is.
There exist exactly four minimal Taylor algebras on a two-element set, counted up to term-equivalence: the two semilattices, the majority algebra, and the affine Mal’cev algebra. Indeed, from the description of their term operations given in Subsection 2.2 it follows that clones of these algebras are even minimal, in the sense that the only proper subclone is the clone of projections. A nice example of a minimal Taylor algebra on a three-element domain is the rock-paper-scissors algebra mentioned in Subsection 3.3. To see that this algebra is minimal Taylor observe that any term operation behaves on any two-element set like the term operation of a two-element semilattice with the same set of essential coordinates. Therefore, the original operation can be obtained by identifying variables in any term operation having at least two essential coordinates. The same argument shows that any semilattice, not necessarily two-element, is minimal Taylor. Affine Mal’cev algebras are also minimal Taylor; this can be deduced e.g. from Theorem 2.4.
In Subsection 5.1 we give the basic general theorems that were proved in [Brady20:bw] in the context of minimal bounded width algebras. Subsection 5.2 concentrates on absorption and related concepts in Zhuk’s theory. It turns out that 2-absorbing sets are exactly projective subuniverses (Theorem 5.7) and 3-absorbing sets are exactly centers (Theorem 5.10). Subsection 5.3 shows that edges substantially simplify in minimal Taylor algebras (Theorem 5.13) and gives additional information for minimal edges; in particular, minimal semilattice edges coincide with thin semilattice edges as defined in [Bulatov16:graph, Bulatov20:local1] (Proposition 5.15). Finally, in Subsection 5.4, we demonstrate a strong interaction between absorption and edges. We show that 2-absorbing subuniverses are exactly subsets that are, in some sense, stable under all the edges (Theorem 5.19), we provide somewhat weaker interaction between absorbing subuniverses and subsets stable under semilattice and abelian edges (Theorem 5.21), we give a common witnessing operation for all of the edges as well as all of the 2- and 3-absorbing subuniverses (Theorem 5.23), and we show that each such a witnessing operation generates the whole clone of term operations (Theorem 5.24). Proofs for this section, including verification of examples, are in Section 9.
@@ 5.1 General facts
Another simple, but important consequence of cyclic operations is the following proposition. The result is slightly more technical than most of the others, but it is in the core of many strong properties of minimal Taylor algebras. The proposition can be further strengthened, see [brady:notes, Theorem 4.2.4.].
@@
(link to proof) Let be a minimal Taylor algebra and $B\subseteq A$ be closed under an operation $f\in\operatorname{Clo}($ such that $B$ together with the restriction of $f$ to $B$ forms a Taylor algebra. Then $B$ is a subuniverse of .
@@ 5.2 Absorption
(link to proof) Let be a minimal Taylor algebra and $B$ an absorbing set of . Then $B$ is a subuniverse of .
@@
Let be an algebra and $B\subseteq A$. The set $B$ is a strongly projective subuniverse of if for every $f\in\operatorname{Clo}_{n}($ and every essential coordinate $i$ of $f$, we have $f({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ is such that $a_{i}\in B$.
@@
The main value of this theorem is the implication showing that, in minimal Taylor algebras, every $2$-absorption, i.e. (a), is as strong as possible (d). Moreover, (b) provides a nice relational description of $2$-absorption, which collapses the general condition from Proposition 2.6 for projectivity to arity $3$. The implications (d) implies (c) implies (b) implies (a) hold for all Taylor algebras. The following examples disprove the reverse implications. We leave the verification as an exercise.
@@
Let $t(x,y,z,w)$ be an operation on $\{0,1\}$ which is equal to $x\vee y\vee z\vee w$ everywhere, except that it is zero on the tuple $(1,0,0,0)$ and its cyclic shifts. Then is a Taylor algebra and the $R$ defined for $B=\{1\}$ in (b) of Theorem 5.7 is compatible with . On the other hand $B$ is not a projective subuniverse of as witnessed by, for example, $t$ and thus (c) of Theorem 5.7 is false.
@@
    If $B\trianglelefteq_{2}$ and $C\leq$, then $B\cup C\leq\mathbf{A}$.
@@
    If $B\trianglelefteq_{2}$ and $C\trianglelefteq$ by $f$, where $\emptyset\neq B,C\neq A$, then
@@
    If , then .
@@
    $\mathbf{A}$ has a unique minimal 2-absorbing subalgebra . Moreover, this algebra does not have any nontrivial 2-absorbing subalgebra.
@@
    there exists with $\operatorname{Clo}($ such that $R(x,y)=B(x)\vee(y=0)$ is a centrality witness.
@@
Just like in Theorem 5.7 we have that a relatively weak notion of $3$-absorption implies a very strong type of centrality which is (d). Let us investigate (d) in greater detail. The fact that $R$ is a subuniverse of translates to the following fact. To every operation of , say $f$, we can associate an operation $f^{\prime}\in\operatorname{Clo}(\{0,1\},\operatorname{maj})$ such that $f({\mathbf{a}})\in B$ whenever $f^{\prime}({\mathbf{x}})=1$ and ${\mathbf{x}}$ is the characteristic tuple of ${\mathbf{a}}$ with respect to $B$ (i.e. $x_{i}=1$ if and only if $a_{i}\in B$). That is, from the viewpoint of “being outside $B$” vs. “being inside $B$” every operation outputs “inside $B$” every time the corresponding operation of $\operatorname{Clo}(\{0,1\};\operatorname{maj})$ outputs $1$.
In fact, there exists a cyclic $t$ in (say, $p$-ary) such that $R(x,y)=B(x)\vee(y=0)$ is a subuniverse of $(A;t)\times(\{0,1\};\operatorname{maj}_{p})$ for every $3$-absorbing $B$. This translates to a simpler statement: for every $3$-absorbing $B$ we have $t({\mathbf{a}})\in B$ whenever majority of the $a_{i}$ belong to $B$, and we cannot expect more, as witnessed by the $2$-element majority algebra. Since $t$ is cyclic it generates the whole clone and, for example, item (2) in Proposition 5.12 below becomes obvious.
@@
(link to verification) Consider the algebra $\mathbf{A}=(\{0,1,2\},m)$ where $m$ is the majority operation such that $m(a,b,c)=a$ whenever |$|\{a,b,c\}|=3$. This algebra is minimal Taylor and the set $C=\{0,1\}$ is an absorbing subuniverse of . However, $C$ is not a center of .
@@
    If $B,C\trianglelefteq_{3}\mathbf{A}$, then $B\cup C\leq$ and $B\cap C\trianglelefteq_{3}$.
@@
    If $\emptyset\neq B,C\trianglelefteq_{3}\mathbf{A}$ and $B\cap C=\emptyset$, then $B^{2}\cup C^{2}$ is a congruence on the subalgebra of with universe $B\cup C$ and the quotient is term-equivalent to a two-element majority algebra.
@@
    If , then .